OpenStudy (anonymous):

how do you solve for x in 2=\frac{-6}{x+4}

OpenStudy (anonymous):

\[2=\frac{-6}{x+4}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

multiply both sides by (x+4)

OpenStudy (ajprincess):

Cross multiply. U will get 2(x+4)=-6 Then remove the brackets. U will get 2*x+2*4=-6 2x+8=-6 Subtract 8 both sides Divide by 2 both sides

OpenStudy (anonymous):

\[2 (x+4 ) = -6\] \[2x + 8 = -6\] \[2x = -14\] \[x = \frac{ -14 }{ 2 } = -7\]

OpenStudy (anonymous):

thank you that helped me out.

OpenStudy (anonymous):

you are welcomed.